# Functional normalizing flow for statistical inverse problems of partial differential equations

### v1

@@ 4.1.1 Functional Householder Flow
The variance function, denoted by $\text{var}_{u}(x)$, can be estimated as follows:
$\text{var}_{u}(x)\approx\frac{1}{m}\sum_{i=1}^{m}(u_{i}(x)-\bar{u}(x))^{2},$
where $x\in\Omega$ is a point residing in the domain $\Omega,\bar{u}$ is the mean function, and $m$ is the sample number. The covariance function can be estimated as follows:
@@
(a) Variance
(b) Covariance
(c) Covariance
Figure 2: The estimated variance and covariance functions obtained by the pCN algorithm (blue solid line), and NF-iVI algorithm (red dashed line). (a): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (b): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+10})\}^{N_{g}-10}_{i=1}$; (c): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+20})\}^{N_{g}-20}_{i=1}$.
(a) NF-iVI
(b) pCN
(c) Difference
@@
Figure 3: The comparison of covariance obtained by two method. (a): The covariance given by NF-iVI method; (b): The covariance given by the pCN method; (c): The difference between the covariance obtained by NF-iVI and pCN sampling methods.
Then we make numerical comparisons of the variance and covariance functions. We show the relative errors between the variance functions and the covariance functions:
@@
where $k$ is an integer, $c_{p}(x,y)$ is the covariance function generated from pCN algorithm, and $c_{N}(x,y)$ is the covariance function generated from NF-iVI method. The relative errors are shown in Table 1. In the table, the notation $\bm{c}$ means the total relative error, and the notation $\{c(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ means the covariance function values on the pair of mesh points $\{(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ ($k=0,10,20$). The numbers below this notation are the relative errors between the vectors obtained by pCN method and NF-iVI, respectively.
@@
Householder Flow $0.1081$ $0.0588$ $0.0981$ $0.1593$
Additionally, we will leverage this problem to demonstrate the discretization invariance of functional Householder flow. To assess the impact of discretization, we discretize the problem using 50, 75, 100, 200, and 300 grid points in a regular mesh. The final results are shown in Figure 4.
(a) Discrete level $n=50$
(b) Discrete level $n=75$
(c) Discrete level $n=100$
diag
(d) Discrete level $n=200$
(e) Discrete level $n=300$
(f) Result of pCN algorithm
Figure 4: The comparison of approximate posterior generated by functional Householder flow with different mesh size $n=\{50,75,100,200,300\}$ respectively. The blue shade area represents the $95\%$ credibility region of approximate posterior.
To further validate the discrete invariance of our model, we computed the $L^{2}$-error between the mean functions of the approximate posterior and the true values across various discrete levels. In Table 2, the numbers above are different discreate level we used for NF-iVI, and the numbers below are the relative errors between the means function obtained by NF-iVI and the truth function, respectively. As detailed in Table 2, the $L^{2}$-errors between the mean functions derived from our algorithm and the true values are consistent across different discretizations. Furthermore, Table 3 presents a comparison of the covariance functions of the approximate posterior for several discrete levels. In the table, the numbers above are different discreate level we used for NF-iVI, and the numbers below are the exact values of the covariance function evaluated at the points $(x_{1},x_{1}),(x_{1},x_{2}),(x_{1},x_{3}),(x_{2},x_{2})$, where $x_{1}=0.3,x_{2}=0.5,x_{3}=0.7$.
Collectively, these findings substantiate the assertion that the approximate posteriors generated by our proposed algorithm are invariant to the different discrete levels.
@@
Discreate Level 50 75 100 200 300
$L^{2}$-Error $0.002697$ $0.002695$ $0.0.02755$ $0.002657$ $0.002756$
@@
Discreate Level 50 75 100 200 300
$c(x_{1},x_{1})$ $0.002886$ $0.003027$ $0.003010$ $0.002754$ $0.002865$
$c(x_{1},x_{2})$ $-0.0010431$ $-0.00107031$ $-0.000904$ $-0.0010427$ $-0.0010700$
$c(x_{1},x_{3})$ $0.000227$ $0.000212$ $0.000250$ $0.000225$ $0.000242$
$c(x_{2},x_{2})$ $0.003005$ $0.002938$ $0.003371$ $0.003377$ $0.003001$
By integrating the numerical findings presented in Figure 4 and Tables 2 and 3 with the theoretical results established in Theorem 2.10, we have successfully demonstrated the discretization invariance of the functional Householder flow.
4.1.2 Functional Projected Transformation Flow
As discussed in Subsection 2.3, functional projected transformation flows are expected to exhibit superior approximation capabilities compared to functional Householder flows. To demonstrate this, we consider the same inverse problem defined in Equation (4.3). In this experiment, we set the dimension of $\text{Im}(\mathcal{F}_{n})$ in the functional projected transformation flow to 20 ($M=20$ in Subsection 2.4). By employing only five layers of functional projected transformation flow, we achieve satisfactory results in approximating the posterior. In contrast, using five layers of functional Householder flow would be insufficient to achieve similar performance.
To construct the flow models, we utilize five layers for the functional projected transformation flow and explore different configurations with 5, 9, 14, 18, and 21 layers for the functional Householder flow. Following Algorithm 1, all the functional normalizing flow models are trained for 5000 iterations ($K=5000$) using 30 samples ($N=30$) in each iteration. The initial learning rate is set to $\alpha_{0}=0.01$, with a step decay learning rate schedule of a multiplicative factor $\tau=0.8$ after every 500 iterations.
Firstly, we compare the estimated mean of the pCN and NF-iVI methods. In subfigure (a) of Figure 5, we present the results obtained from the 5-layer functional projected transformation flow. We observe that these results are consistent with the posterior generated by the pCN algorithm, as shown in subfigure (b) of Figure 1, indicating the effectiveness of our proposed method. In subfigures (b)-(f) of Figure 5, we present the results obtained from the functional Householder flow with 5, 9, 14, 18, and 21 layers, respectively. We observe that when the number of layers is small, the functional Householder flow may not achieve satisfactory approximations.
@@ 4.1.2 Functional Projected Transformation Flow
Figure 5: The comparison of the approximation capabilities of functional Householder flow and functional projected transformation flow. The blue shade area represents the $95\%$ credibility region of approximate posterior. (a): Result of functional projected transformation flow with 5 layers. (b)-(f): Result of functional Householder flow with 5, 9, 14, 18, and 21 layers, respectively.
Secondly, we compare the variance and covariance functions generated by the two methods. For simplicity, we compute these quantities on the mesh points and present the results in Figure 7. In all the subfigures of Figure 7, the estimated posteriors obtained by the pCN method, the 5-layer functional Householder flow, and the 5-layer projected transformation flow are drawn in blue solid line, red dashed line, and purple dashed line, respectively. In subfigures (a) and (d) of Figure 7, we show the variance function calculated on all of the mesh points, i.e., $\{\mathrm{var}_{u}(x_{i})\}_{i=1}^{N_{g}}~{}(N_{g}=101$ is the number of mesh points). In subfigures (b) and (e) of Figure 7, we show the covariance function calculated on the pairs of points $\left\{(x_{i},x_{i+10})\right\}_{i=1}^{N_{g}-10}$, and in subfigures (c) and (f) of Figure 7, we show the covariance function calculated on the pairs of points $\{(x_{i},x_{i+20})\}_{i=1}^{N_{g}-20}$, respectively. We also show the matrix representations of the covariance operators $\mathcal{C}_{\text{N}}$ and $\mathcal{C}_{\text{p}}$ along with the difference in subfigures (a)-(c) of Figure 6. Here, $\mathcal{C}_{N}$ is the covariance operator of the approximate posterior generated by the 5-layer projected transformation flow, and $\mathcal{C}_{p}$ is the covariance operator of the estimated posterior generated by the pCN method. The results confirm that the NF-iVI provides a reliable approximation of the covariance function.
(a) NF-iVI
(b) pCN
(c) Difference
@@
Figure 6: The comparison of posterior covariance of $u$ with the mesh size $n=100$ respectively. (a): The covariance given by NF-iVI method; (b): The covariance given by the pCN sampling method; (c): The difference between the covariance obtained by NF-iVI and pCN sampling methods.
Next, we make numerical comparisons of the variance and covariance functions. We use the 5-layer functional Householder flow and the 5-layer functional projected transformation flow to approximate the posterior, and we show the relative errors of the covariance functions in Table 4. In the table, the notation $\bm{c}$ means the total relative error, and the notation $\{c(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ means the covariance function values on the pair of mesh points $\{(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ ($k=0,10,20$). The numbers below this notation are the relative errors between the vectors obtained by pCN method and two different NF-iVI methods, respectively.
Householder Flow $1588.736$ $630.159$ $12867.031$ $596.937$
Projected Transformation $0.03057$ $0.00390$ $0.12691$ $0.03290$
As depicted in Table 4, the relative errors of the 5-layer projected transformation flow are consistently smaller than those of the Householder flow. The variance and covariance functions obtained by the pCN method and the projected transformation flow demonstrate quantitative similarity. This indicates that the 5-layer projected transformation flow exhibits superior performance in approximating the covariance function of the posterior compared to the 5-layer Householder flow.
(a) Variance
(b) Covariance
(c) Covariance
(d) Variance
(e) Covariance
(f) Covariance
Figure 7: The estimated variance and covariance functions obtained by the pCN method (blue solid line), Householder NF-iVI method (green dashed line) and projected transformation NF-iVI (red dashed line). (a)(d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (b)(e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+10})\}^{N_{g}-10}_{i=1}$; (c)(f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+20})\}^{N_{g}-20}_{i=1}$.
The comparisons presented above highlight the superior approximation capabilities of the functional projected transformation flow, even when using a small number of layers. To demonstrate the discrete invariance of the functional projected transformation flow, we conduct a similar analysis to that performed for the functional Householder flow. We calculate the $L^{2}$-errors between the mean of the approximate posterior and the true value under different discrete levels. In Table 5, the numbers above are different discreate level we used for NF-iVI, and the numbers below are the relative errors between the means function obtained by NF-iVI and the truth function, respectively. As detailed in Table 5, the $L^{2}$-errors between the mean functions derived from our algorithm and the true values are consistent across different discretizations. As presented in the table, the $L^{2}$-errors between the mean of the approximate posterior and the true value remains consistent across different discrete levels.
Furthermore, Table 6 presents a comparison of the covariance functions of the approximate posterior for several discrete levels. In the table, the numbers above are different discreate level we used for NF-iVI, and the numbers below are the exact values of the covariance function evaluated at the points $(x_{1},x_{1}),(x_{1},x_{2}),(x_{1},x_{3}),(x_{2},x_{2})$, where $x_{1}=0.3,x_{2}=0.5,x_{3}=0.7$.
Discreate Level 50 75 100 200 300
$L^{2}$-Error $0.003156$ $0.003027$ $0.003117$ $0.003182$ $0.003041$
Discreate Level 50 75 100 200 300
$c(x_{1},x_{1})$ $0.003670$ $0.003993$ $0.003926$ $0.003626$ $0.003948$
$c(x_{1},x_{2})$ $-0.001043$ $-0.001158$ $-0.001165$ $-0.001062$ $-0.001011$
$c(x_{1},x_{3})$ $0.000289$ $0.000349$ $0.000273$ $0.000207$ $0.000251$
$c(x_{2},x_{2})$ $0.003414$ $0.003521$ $0.003545$ $0.003805$ $0.003817$
@@ 4.2 Steady-State Darcy Flow Equation
In this subsection, we focus on the inverse problem of estimating the permeability distribution in a porous medium from a discrete set of pressure measurements, as studied in [11]. Consider the following steady-state Darcy flow equation:
@@
where $f\in H^{-1}(\Omega)$ is the source function, and $u\in\mathcal{H}_{u}:=L^{\infty}(\Omega)$ is called log-permeability for the computational area $\Omega=(0,1)^{2}$. The forward operator has the following form:
$\mathcal{S}\mathcal{G}u=(w(x^{1}),w(x^{2}),\ldots,w(x^{N_{d}}))^{T},$
@@
$\bm{d}=\mathcal{S}\mathcal{G}u+\bm{\epsilon},$ (4.7)
@@
    Let the domain $\Omega$ be a bounded area $(0,1)^{2}$. The measurement points $\{x^{i}\}^{N_{d}}_{i=1}$ are taken at the coordinates $\{\frac{i}{21},\frac{j}{21}\}_{i,j=1}^{20}$.
@@ 4.2.1 Functional Planar Flow
Given the nonlinear nature of the inverse problem, we employ a functional planar flow model to approximate the posterior distribution, constructing it with 24 layers. Following Algorithm 1, the functional planar flow is trained over 5000 iterations $(K=5000)$, with each iteration utilizing 30 samples $(N=30)$. We utilize a step decay learning rate schedule, starting with an initial learning rate of $\alpha_{0}=0.01$ and decreasing it by a multiplicative factor of $\tau=0.8$ after every 500 iterations. For comparison, we also present the posterior generated by the pCN algorithm. In the discussion of all steady-state darcy flow problems, when we discuss ESS, the corresponding functional $f$ is defined as $f(u)=u((0.5,0.5))$. We generate $3\times 10^{6}$ samples using the pCN algorithm, achieving an effective sample size of $\text{ESS}_{p}=278$.
@@
Firstly, in Figure 8, we present a comparative analysis of the mean function generated by the pCN algorithm and the approximate posterior obtained from functional planar flow. We observe that the mean of the approximate posterior closely resembles the ground truth, as shown in subfigures (a) and (b). Additionally, subfigures (d) and (e) of Figure 8 illustrate the mean function of the estimated posterior generated by the pCN algorithm. The similarity between these results further demonstrates the effectiveness of our proposed method in accurately estimating the mean function.
@@
(b) Estimated mean
(c) Estimated variance
@@
(e) Estimated mean
(f) Estimated variance

### v2

@@ 4.1 Simple Elliptic Equation
The variance function, denoted by $\text{var}_{u}(x)\mathrel{\mathop{\ordinarycolon}}=c(x,x)$, can be estimated as
$\text{var}_{u}(x)\approx\frac{1}{m-1}\sum_{i=1}^{m}(u_{i}(x)-\bar{u}(x))^{2},$
where $x\in\Omega$ is a point residing in the domain $\Omega,\bar{u}$ is the mean function, and $m$ is the number of samples. The covariance function can be estimated as
@@
(a) variance (Householder)
(b) covariance (Householder)
(c) covariance (Householder)
(d) variance (projected)
(e) covariance (projected)
(f) covariance (projected)
Figure 2: The estimated variance and covariance functions obtained by the pCN algorithm (blue solid line), and NF-iVI algorithm (red dashed line). (a)(d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (b)(e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+10})\}^{N_{g}-10}_{i=1}$; (c)(f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+20})\}^{N_{g}-20}_{i=1}$.
(a) covariance (Householder)
(b) covariance (projected)
(c) covariance (pCN)
@@
Figure 3: The comparison of covariance. (a): The covariance given by Householder NF-iVI method; (b): The covariance given by Projected NF-iVI method; (c): The covariance given by the pCN method;
Then we make numerical comparisons of the variance and covariance functions by showing the relative errors:
@@
where $k$ is an integer, $c_{p}(x,y)$ is the covariance function generated from pCN algorithm, and $c_{N}(x,y)$ is the covariance function generated from NF-iVI method. The relative errors are shown in Table 1. Here, the notation $\bm{c}$ means the total relative error, and the notation $\{c(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ means the covariance function values on the pair of mesh points $\{(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ ($k=0,10,20$). The numbers below this notation are the relative errors between the vectors obtained by pCN method and NF-iVI, respectively.
@@
Householder Flow $0.03513$ $0.01911$ $0.09921$ $0.00921$
Projected Transformation $0.03656$ $0.00686$ $0.11220$ $0.03715$
Additionally, we will leverage this problem to demonstrate the discretization invariance of our flow model. To maintain conciseness, we present the discretization-invariance numerical experiments only for the Householder flow. The analogous results for the other three proposed flow models, which exhibit similar properties, are omitted but can be fully reproduced and verified using the publicly available code.
To assess the impact of discretization, we discretize the problem using 50, 75, 100, 200, and 300 grid points in a regular mesh, and compute the $L^{2}$-error between the mean functions of the approximate posterior and the background truth across various discrete levels. In Table 2, the numbers above are different discrete level we used for NF-iVI, and the numbers below are the relative errors between the means function obtained by NF-iVI and the true function, respectively. As detailed in Table 2, the $L^{2}$-errors between the mean functions derived from our algorithm and the truth are consistent across different discretizations. Furthermore, Table 3 presents a comparison of the covariance functions of the approximate posterior for several discrete levels. In the table, the numbers above are different discrete level we used for NF-iVI, and the numbers below are the exact values of the covariance function evaluated at the points $(x_{1},x_{1}),(x_{1},x_{2}),(x_{1},x_{3}),(x_{2},x_{2})$, where $x_{1}=0.3,x_{2}=0.5,x_{3}=0.7$. Collectively, these findings substantiate the assertion that the approximate posteriors generated by our proposed algorithm are invariant to the different discrete levels.
@@
Discrete Level 50 75 100 200 300
$L^{2}$-Error $0.00286$ $0.00294$ $0.00305$ $0.00311$ $0.00302$
@@
Discrete Level 50 75 100 200 300
$c(x_{1},x_{1})$ $0.00353$ $0.00340$ $0.00323$ $0.00326$ $0.00315$
$c(x_{1},x_{2})$ 0.00120-0.00120 0.00123-0.00123 0.00122-0.00122 0.00130-0.00130 0.00127-0.00127
$c(x_{1},x_{3})$ $0.00028$ $0.00036$ $0.00016$ $0.00030$ $0.00023$
$c(x_{2},x_{2})$ $0.00299$ $0.00272$ $0.00290$ $0.00329$ $0.00287$
By integrating the numerical findings presented in Tables 2 and 3 with the theoretical results established in Theorem 2.9, we have successfully demonstrated the discretization invariance of the flow model.
As discussed in Subsection 2.4, functional projected transformation flow are expected to exhibit superior approximation capabilities compared with functional Householder flow. To demonstrate this, we utilize 5 layers for the functional projected transformation flow and explore different configurations with 5, 9, 14, 18, and 21 layers for the functional Householder flow. Following Algorithm 1, all the functional normalizing flow models are trained for 5000 iterations ($K=5000$) using 30 samples ($N=30$) in each iteration. The initial learning rate is set to $\alpha_{0}=0.01$, with a step decay learning rate schedule of a multiplicative factor $\tau=0.8$ after every 500 iterations.
In subfigure (a) of Figure 4, we present the results obtained from the 5-layer functional projected transformation flow. We observe that these results are consistent with the posterior generated by the pCN algorithm, as shown in subfigure (c) of Figure 1, indicating the effectiveness of our proposed method. In subfigures (b)-(f) of Figure 4, we present the results obtained from the functional Householder flow with 5, 9, 14, 18, and 21 layers, respectively. We observe that when the number of layers is small, the functional Householder flow may not achieve satisfactory approximations.
@@
Figure 4: The comparison of the approximation capabilities of functional Householder flow and functional projected transformation flow. The green shade area represents the $95\%$ credibility region of approximate posterior. (a): Result of functional projected transformation flow with 5 layers. (b)-(f): Result of functional Householder flow with 5, 9, 14, 18, and 21 layers, respectively.
The comparisons presented above highlight the superior approximation capabilities of the functional projected transformation flow, even when using a small number of layers. Furthermore, to demonstrate the advantage of the Bayesian approach in solving inverse problems, we utilize a reduced number of measurement points to investigate the resulting posterior distribution under the constraint of sparse measurement data. Our algorithm employ a 5-layer projected transformation flow. The inversion is performed under three measurement scenarios, described below:
  • 1.
    Full measurement: The measurement data are assumed to be $\{w(x^{i})|x^{i}=i/10,i=0,1,2,\ldots,10\}$.
  • 2.
    Left measurement: The measurement data are assumed to be $\{w(x^{i})|x^{i}=i/10,i=0,1,\ldots,5\}$.
  • 3.
    Right measurement: The measurement data are assumed to be $\{w(x^{i})|x^{i}=i/10,i=5,6,\ldots,10\}$.
Full measurement scenario involves measurement points spread across the domain $\Omega$. Left measurement scenario localizes measurement points on the left side of $\Omega$, and right measurement scenario localizes the measurement points on the right. All other settings for the inverse problem remain unchanged. The resulting posterior corresponding to these three distinct measurements are presented in Figure 5 below.
(a) full measurement
(b) left measurement
(c) right measurement
@@
Figure 5: A comparison of the approximate posterior with different measurement scenarios. The green shade area represents the $95\%$ credibility region of estimated posterior. (a): Posterior of full measurement. (b): Posterior of left measurement. (c): Posterior of right measurement.
Subfigures (a), (b), and (c) of Figure 5 present the results of the inverse problem with full measurement, left measurement, and right measurement, respectively. It is evident that for this linear inverse problem, the posterior distribution—when informed by full measurements—accurately reflects the true state of the model parameters. Conversely, when observations are sparse (specifically, when measurements are concentrated on the left side of the domain while the right side is neglected), the posterior distribution exhibits narrow credible intervals on the left, with the posterior mean closely matching the truth. In sharp contrast, the right side is characterized by significantly wider credible intervals, and the posterior mean deviates substantially from the ground truth. An analogous behavior is observed when measurements are concentrated on the right side of the domain.
These results do not reflect a deficiency in the model’s performance under sparse measurement settings; instead, they indicate that the measurement data are insufficient to accurately recover the true model parameters in both left-side and right-side measurement scenarios. When measurement data are inadequate, the credible regions provided by the Bayesian posterior are critical, as they delineate the regions where accurate inference is feasible and, conversely, where reliable inference cannot be achieved.
@@ 4.2 Steady-State Darcy Flow Equation
In this subsection, we focus on the inverse problem of estimating the permeability distribution in a porous medium from a discrete set of pressure measurements, as studied in calvetti2018iterative. Consider the following steady-state Darcy flow equation:
@@
where $f\in H^{-1}(\Omega)$ is the source function ($f=sin(\pi x_{1})sin(\pi x_{2})$ in our experiment, where $x=(x_{1},x_{2})$), and $u\in\mathcal{H}_{u}\mathrel{\mathop{\ordinarycolon}}=L^{\infty}(\Omega)$ is called log-permeability for the computational area $\Omega=(0,1)^{2}$. The forward operator has the following form:
$\mathcal{S}\mathcal{G}(u)=(w(x^{1}),w(x^{2}),\ldots,w(x^{N_{d}}))^{T},$
@@
$\bm{d}=\mathcal{S}\mathcal{G}(u)+\bm{\epsilon},$ (4.8)
@@
    Let the domain $\Omega$ be a bounded area $(0,1)^{2}$. The measurement points $\{x^{i}\}^{N_{d}}_{i=1}$ are taken at the coordinates $\{i/21,j/21\}_{i,j=1}^{20}$.
@@
Given the nonlinear nature of the inverse problem, we employ two nonlinear flows, i.e., functional planar flow and functional Sylvester flow to approximate the posterior. Functional planar flow is constructed with 32 layers, while functional Sylvester flow is constructed with 5 layers. Following Algorithm 1, all the flow models are trained over 5000 iterations $(K=5000)$, with each iteration utilizing 30 samples $(N=30)$. We utilize a step decay learning rate schedule, starting with an initial learning rate of $\alpha_{0}=0.01$ and decreasing it by a multiplicative factor of $\tau=0.8$ after every 500 iterations. For comparison, we also present the posterior generated by the pCN algorithm with $3\times 10^{6}$ samples. It is worth noting that running the pCN algorithm required approximately 472 minutes, whereas training the FNF model took only about 28 minutes.
@@
Firstly, in Figure 6, we present a comparative analysis of the mean function generated by the pCN algorithm and the approximate posterior obtained from the flow models. Subfigures (a), (b), and (c) of Figure 6 illustrate the mean function of the estimated posterior. The similarity between these results demonstrates the effectiveness of our proposed method in accurately estimating the mean function.
(a) mean (pCN)
(b) mean (planar)
(c) mean (Sylvester)
diag
(d) variance (pCN)
(e) variance (planar)
(f) variance (Sylvester)
Figure 6: The comparison of the measures obtained by flow models and pCN algorithm. (a)(b)(c): The mean function of posterior obtained by pCN algorithm, 32 layers planar flow, 5 layers Sylvester flow, respectively. (d)(e)(f): The variance function of posterior obtained by pCN algorithm, 32 layers planar flow, 5 layers Sylvester flow, respectively.
To further support the effectiveness of the NF-iVI method in estimating the mean function, we provide numerical evidence. The relative errors between the mean of the approximate posterior generated by 32 layers planar flow, 5 layers Sylvester flow and the mean of the posterior $u_{p}^{*}$ generated by the pCN algorithm are given by:
$\textbf{planar relative error:}=\frac{\lVert u_{pN}^{*}-u_{p}^{*}\rVert^{2}}{\lVert u_{p}^{*}\rVert^{2}}=0.00142,$ (4.9)
$\textbf{Sylvester relative error:}=\frac{\lVert u_{SN}^{*}-u_{p}^{*}\rVert^{2}}{\lVert u_{p}^{*}\rVert^{2}}=0.00188,$ (4.10)
where $u_{SN}^{*}$ is the mean of planar flow, and $u_{SN}^{*}$ is the mean of Sylvester flow. The estimated values are small, indicating the effectiveness of our algorithm.
Combining the visual evidence from Figure 6 and the numerical evidence provided by the relative error in Equations (4.9) and (4.10), we conclude that the NF-iVI method effectively approximates the mean of the posterior.
Secondly, to further explore the approximate posterior, we analyze the covariance functions. In subfigures (e)(f) of Figure 6, we present the point-wise variance field of the posterior measure obtained from the NF-iVI method. For comparison, subfigure (d) of Figure 6 depicts the point-wise variance field estimated from pCN method. The similarity between these three plots indicates the effectiveness of the NF-iVI method in quantifying the uncertainties of the parameter $u$.
We also compare the covariance functions generated by these two methods. For simplicity, we compute these quantities at the mesh points and present the results in Figure 7. In all subfigures of Figure 7, the estimated covariance of posteriors obtained by the pCN and NF-iVI methods are represented by a blue solid line and a red dashed line, respectively. In subfigures (a)(d) of Figure 7, we show the variance function calculated at all mesh points, i.e., $\{\mathrm{var}_{u}(x_{i})\}_{i=1}^{N_{g}}~(N_{g}$ is the number of mesh points, which is 441 in this problem). In subfigures (b)(e) and (c)(f) of Figure 7, we show the covariance functions calculated on the pairs of points $\left\{(x_{i},x_{i+40})\right\}_{i=1}^{N_{g}-40}$ and $\{(x_{i},x_{i+80})\}_{i=1}^{N_{g}-80}$, respectively.
To numerically compare the variance and covariance functions generated by the two methods, we present the relative errors in Table 4. In the table, the notation $\bm{c}$ means the total relative error, and the notation $\{c(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ means the covariance function values on the pair of mesh points $\{(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ ($k=0,40,80$). The numbers below this notation are the relative errors between the vectors obtained by pCN method and NF-iVI, respectively. The small relative errors indicate that the posterior covariance functions obtained by the NF-iVI and pCN sampling methods are quantitatively similar. Combining the visual evidence from Figures 6 and 7, and the numerical evidence from Table 4, we conclude that the NF-iVI method effectively approximates the covariance function of the posterior.
Planar Flow $0.07252$ $0.00889$ $0.01254$ $0.03345$
Sylvester Flow $0.09725$ $0.01227$ $0.02312$ $0.05118$
(a) variance (planar)
(b) covariance (planar)
(c) covariance (planar)
(d) variance (Sylvester)
(e) covariance (Sylvester)
(f) covariance (Sylvester)
Figure 7: The estimated variance and covariance functions obtained by the pCN algorithm (blue solid line), and NF-iVI algorithm (red dashed line). (a)(d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (b)(e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$; (c)(f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$.
As discussed in Subsection 2.5, the functional Sylvester flow is anticipated to exhibit superior approximation capabilities compared with the functional planar flow. To demonstrate this, we consider the same inverse problem defined in equation (4.8). In this experiment, we set the dimension of $\text{Im}(\mathcal{F}_{n})$ in functional Sylvester flow to 20 ($M=20$ in Subsection 2.5). By employing only 5 layers of functional Sylvester flow, we achieve satisfactory results in approximating the posterior. In contrast, using 5 layers of functional planar flow would be insufficient to achieve similar performance.
Subfigures (a) and (d) of Figure 8 show the ground truth. Subfigures (b), (c), (e), and (f) depict the mean functions and the estimated pointwise variance fields of the approximate posterior obtained using the 5-layer functional Sylvester flow and the 5-layer functional planar flow. The comparisons highlight the limitations of functional planar flow in achieving accurate approximations with a small number of layers. In contrast, functional Sylvester flow exhibits superior approximation capabilities.
@@
(b) mean (Sylvester-5)
(c) variance (Sylvester-5)
@@
(e) mean (planar-5)
(f) variance (planar-5)